A donut crate has 100 apple fritters, 150 chocolate fritters, 200 apple bars and 250 chocolate bars. Choose one at random. What is the probability it is a fritter or contains chocolate?
1
Expert's answer
2013-02-18T08:39:15-0500
The number of fritter or containing chocolate is 100+150+250=500.
The total number of items is 100+150+200+250=700.
The probability of choosing a fritter or containing chocolate is: P=500/700=5/7 or approximately 71.4%